# LAP, Using Action Feasibility for Improved Uncertainty Alignment of Large Language Model Planners

### v2

@@ Abstract
Large language models (LLMs) showcase many desirable traits for intelligent and helpful robots. However, they are also known to hallucinate predictions. This issue is exacerbated in robotics where LLM hallucinations may result in robots confidently executing plans that are contrary to user goals, relying more frequently on human assistance, or preventing the robot from asking for help at all. In this work, we present LAP, a novel approach for utilizing off-the-shelf LLMs, alongside a novel Action feasibility metric, in robotic Planners that minimize harmful hallucinations and human intervention. Our key finding is that calculating and leveraging a new metric, which we call A-Feasibility, a measure of whether a given action is possible and safe in the provided scene, helps to mitigate hallucinations in LLM predictions and better align the LLM’s confidence measure with the probability of success. We specifically propose an A-Feasibility metric which both combines scene context and prompting a LLM to determine if a given action is possible and safe in the scene, using the LLM’s response to compute the score. Through experiments in both simulation and the real world on tasks with a variety of ambiguities, we show that LAP significantly increases success rate and decreases the amount of human intervention required relative to prior art. For example, in our real-world testing paradigm, LAP decreases the human help rate of previous methods by over 33% at a success rate of 70%.
@@ I Introduction
Figure 1: LLMs are not grounded in the real world and will frequently hallucinate. Additionally, when provided with an ambiguous instruction, the LLM must know when to ask users for help. We present LAP, a novel approach for calibrating LLM confidence with an action feasibility metric to better detect hallucinations and resolve ambiguities, before asking a user for help when necessary. We calculate a novel metric, which we call A-Feasibility, for each possible action and combine it with the next token probability of that action to determine both which option is most likely, and if help is needed. Note, in the provided example, ‘pick up coffee cup’ is a hallucination caused by the users phrasing of their command. No such specific object was grounded in the provided ground truth perception information. In LAP, our A-Feasibility metric, seen in orange, mitigate this case.
@@
Main Contributions: We introduce LAP, a novel approach for calibrating LLM confidence with action feasibility. Our key finding is that a new metric, which we call A-Feasibility, which can intuitively be thought of as the potential for an action to be both possible and safe in the given scene, can minimize hallucinations, improve safety, and more accurately calibrate uncertainty. By handling these issues, the A-Feasibility metric can be used to increase the task success rate and minimize the frequency of human intervention. An additional advantage of our A-Feasibility metric is that it does not require extensive training, which is difficult to do without overfitting for extremely diverse and data-limited tasks like those found in home robotics.
@@
    We introduce LAP, a novel approach for uncertainty alignment that uses Large Language Models and a feasibility to better align model confidence with task success for high-level robotic planners.
@@
    We evaluate our approach in both simulation and on hardware in the real world using a suite of language-instructed manipulation tasks from the KnowNo Simulation and Mobile Manipulator Datasets [3]. We show that LAP significantly increases the success rate and reduces the amount of help needed as compared to our baselines across different environments and LLMs. Specifically in our real-world testing, we see a decrease in the human help rate of previous methods by over 33% at a success rate of 70%, with a similar decrease at most success rates.
@@
    We explore and evaluate different ways of extracting the A-Feasibility metric for use in LAP. We propose the use of a combined metric which utilizes scene context alongside a novel prompt-based metric, which allows added flexibility in the A-Feasibility metric to encapsulate safety and other important behavior.
@@
    We show that using LAP on the newest, most powerful language models out-of-the-box outperforms fine-tuning LAP and prior art [3] on the best models available for fine-tuning on this task.
Figure 2: We present LAP, a novel approach for uncertainty alignment that utilizes a novel A-Feasibility metric. After generating a set of multiple-choice candidate actions, each prepended with ‘A’, ‘B’, ‘C’, etc., we calculate the A-Feasibility metric for each option (shown in orange). One version of our method uses the scene context from earlier prompts (shown in blue), while another sends a new query to the LLM to score a given option. This prompt-based method allows us to include factors like safety into the A-Feasibility metric. After computing each option’s A-Feasibility metric, we multiply the next-token likelihood and construct a prediction set of options with a total value above a threshold, $t$. Note that we eliminate option A, ”Put metal bowl in the microwave,” as a valid option as it is dangerous. If the size of the prediction set were greater than one, the agent would ask for help.
@@ II-A Hallucinations and Uncertainty in LLMs
As LLMs have risen to the fore, an increasing amount of work has been aimed at addressing their hallucinations, times when the LLM generates content not grounded in reality. Some works aim to quantify the frequency or severity of hallucinations [4] in a given LLM, while many aim to decrease their frequency [5], or better calibrate uncertainty [9, 10, 11]. Some recent works attempt to handle hallucinations by using conformal prediction-based methods to provide coverage guarantees[12]. Ren et al. [3] introduce KnowNo, which uses conformal prediction to create coverage guarantees for robotics tasks, asking users for help when the agent is unsure. In contrast to these works, we mitigate hallucinations through the introduction of our A-Feasibility metric, minimizing the amount of help needed.
II-B Feasibility and Robotics
@@
We take inspiration from the use of affordances and extend its definition to what we call action feasibility, or the potential to complete an action in the given scene safely and effectively. We also computer our A-Feasibility metric without the need for training or LLM fine-tuning. We choose to do this as training capable value functions for every possible action in the real world (as SayCan does) is extremely difficult due to a lack of domain data and vast task diversity. We also show that using feasibility can help mitigate hallucinations in an LLM’s generations by grounding the generations in the real world.
@@ II-C LLMs for Robot Planning and Human-Robot Interaction
III A-Feasibility Metric for Decreasing Uncertainty
@@
In this work, we consider the following problem statement: Given a robotic agent and a user-provided instruction, execute the instruction completely and accurately while resolving any ambiguities as needed through dialogue with the user. Minimize the amount of dialogue. Specifically, our robot will begin in a set environment, awaiting a user’s instruction. The user will provide an instruction, which may be under-specified, at which point our method will need to evaluate whether or not it requires additional information from the user. This problem space was initially explored by [3]. Following [3], we choose to transform the problem into an multiple choice question answer (MCQA) task, which allows us to define the problem as next-token prediction – aligning well with LLM loss functions and training data. To do this, an LLM is provided a few-shot prompt (provided in supplementary materials) that includes possible next steps in a few example scenarios. The LLM in turn generates a set, $\{y^{i}\}$, of candidate actions that are semantically different for the current scenario. Then, the task of choosing among the candidate actions is formatted as a multiple choice question to the LLM, with each option preceded with ‘A’, ‘B’, ‘C’, ‘D’, or ‘E’ as shown in Figure 2. The probabilities of each possible token (i.e. the multiple choice options ‘A’, ‘B’, ‘C’, …) then serve as normalized scores that can be utilized by uncertainty quantification methods, like LAP and KnowNo [3]. We utilize these normalized scores alongside our A-Feasibility metric to generate a prediction set of plans from $\{y^{i}\}$. Please see [3] for further rationale on why MCQA should be used for uncertain robot planning.
Method Overview. We present an overview of our method, LAP, in Figure 2. LAP begins with the MCQA paradigm described above, outputting the model’s uncalibrated confidence score for each generated action option, $p^{*}(y^{i})$. We then compute the A-Feasibility metric, $a^{i}$, for each option, corresponding to the plausibility of that action given the scene. We describe three different methods of obtaining this score below, context-based, perception-based, and prompt-based. Each of these scores can be used independently or combined into a superscore as shown in Figure 2. We choose to present all three as each could be advantageous in different use cases and their combination can create a more performative final score. We now define our final probability of each option being the user’s intent as:
$p(y^{i})=p^{*}(y^{i})\cdot a^{i}.$ (1)
@@
The LLM is certain if this set is a singleton, and queries the user for help if not. Calculating an optimal $t$ is the main subject of KnowNo [3], which uses conformal prediction to provide statistical guarantees of a certain success rate. Our work differs significantly from KnowNo as we improve $p(y^{i})$ by integrating our A-Feasibility metric.
Context-Based A-Feasibility Metric. Our first method of calculating the A-Feasibility metric that we explore is the most efficient, requiring the least amount of time, compute, and control over the robot agent. This method uses the context of the scene that the robot already found and provided to the LLM from its perception algorithms. We label this A-Feasibility metric $a^{i}_{c}$. While perception information is provided, this does not always prevent the LLM from hallucinating new objects into existence.These types of hallucinations become even more prominent when the hallucinated object is in the user command. An example of this can be seen in Figure 1 with the ‘coffee mug’ object.
@@
To calculate $a^{i}_{c}$, the scene context is first parsed into a set of found objects, $\mathcal{O}$. Each generated action candidate from the LLM’s MCQA creation response, $y^{i}$, is parsed for any objects, $o^{i}$, present. If all of the objects in $y^{i}$ are in the set of found objects, $a^{i}_{c}=1$. Otherwise, $a^{i}_{c}=0$. To formalize this:
$a^{i}_{c}=\begin{cases}1&\text{if }\forall o^{i}\in y^{i},\,o^{i}\in\mathcal{O}\\ 0&\text{otherwise}.\end{cases}$ (3)
@@
Perception-Based A-Feasibility Metric. Building off of the context-based A-Feasibility metric, the perception-based A-Feasibility metric takes into account any uncertainty in the perception algorithms themselves. To do this, we run an open-vocabulary vision-language model, like CLIP [33] or ViLD [34], on the scene observations with each object, $o^{i}$ (from the option generated by the LLM, $y^{i}$) as input.
@@
Specifically, we take the camera images of the scene, $s$, and feed them through ViLD with the list of objects $\{o^{i}\}$ as input. We then take the probabilities of each object $p(o^{i}|s)$ as outputted by the model and find the average for a given option, $y^{i}$. Formally, we define $a^{i}_{perc}$ as
$a^{i}_{perc}=\frac{1}{n}\sum_{i=1}^{n}p(o^{i}|s),$ (4)
@@
for $n$ objects in $y^{i}$. If the object is localized to the same location as an object in the scene context using IoU (usually an indicator of a hallucination), the metric is lowered to 0 to prevent duplicate options of the same object.
@@
Prompt-Based A-Feasibility Metric. Recent works like [5] and [22] have shown that additional prompts to the LLM requiring it to reason on its prior outputs can isolate harmful hallucinations. We leverage this idea and propose a prompt-based A-Feasibility metric for each action option, $y^{i}$, generated by the model. This additionally allows us to generate a more flexible A-Feasibility metric, which can account for factors beyond just the grounding of objects in the scene. Our prompt, given our real-world evaluation environment, begins with a series of few-shot samples, before serving the following text:
@@
In the prompt, the items inside the brace notation are replaced with their respective values. The header includes the description of items found in the scene from the MCQA generation and scoring prompts and describes the task for the LLM: determining whether the given action is ‘possible and safe’ in the current scene. The header also provides a few examples. Once we receive the response from the LLM, we extract the normalized probability of the ‘True’ token and utilize this as our A-Feasibility metric $a^{i}_{pr}$, for the option $y^{i}$.
Figure 3: Comparison of task success rate vs human help rate (left) and average prediction set size (right) in our Simulation experiments. Note LAP improving upon KnowNo for almost all success rates. Including a reduction in the human help rate of over 70% at a success rate of 60%. No Help and Binary do not create prediction sets and are not included in the plot.
One benefit of this approach is the ability to tailor the A-Feasibility metric to new criteria. For example, in this prompt, we add safety as a major component of the metric, allowing the LLM to lower the metric for putting the metal bowl in the microwave. Similarly, if an important characteristic of the robot agent is that it does not leave a certain area, one could specify this and the A-Feasibility metric should adapt. The perception-based methods above cannot apply logic to the generated options, $y^{i}$, in this way.
Fine-Tuning. We also hypothesize that utilizing the newest, most powerful, LLM available off-the-shelf will generally outperform fine-tuning the older models available for fine-tuning. We tested this hypothesis by fine-tuning gpt-3.5-turbo-1106 [35], the most powerful model available from OpenAI for fine-tuning. For the simulation environment, we fine-tuned on each of the three LLM phases of LAP, the MCQA generation, MCQA scoring, and A-Feasibility scoring.

### v3

# LBAP: Improved Uncertainty Alignment of LLM Planners using Bayesian Inference

@@ Abstract
Large language models (LLMs) showcase many desirable traits for intelligent and helpful robots. However, they are also known to hallucinate predictions. This issue is exacerbated in robotics where LLM hallucinations may result in robots confidently executing plans that are contrary to user goals or relying more frequently on human assistance. In this work, we present LBAP, a novel approach for utilizing off-the-shelf LLMs, alongside Bayesian inference for uncertainty Alignment in robotic Planners that minimizes hallucinations and human intervention. Our key finding is that we can use Bayesian inference to more accurately calibrate a robots confidence measure through accounting for both scene grounding and world knowledge. This process allows us to mitigate hallucinations and better align the LLM’s confidence measure with the probability of success. Through experiments in both simulation and the real world on tasks with a variety of ambiguities, we show that LBAP significantly increases success rate and decreases the amount of human intervention required relative to prior art. For example, in our real-world testing paradigm, LBAP decreases the human help rate of previous methods by over 33% at a success rate of 70%.
@@ I Introduction
Figure 1: LLMs are not grounded in the real world and will frequently hallucinate. Additionally, when provided with an ambiguous instruction, the LLM must know when to ask users for help. We present LBAP, a novel approach for calibrating LLM confidence with Bayesian inference to better detect hallucinations and resolve ambiguities, before asking a user for help when necessary. Note, in the provided example, ‘pick up coffee cup’ is a hallucination caused by the users phrasing of their command. No such specific object was grounded in the perception information. In LBAP, Bayesian inference incorporates this fact into the final probabilities, allowing us to mitigate this case.
@@
Main Contributions: We introduce LBAP, a novel approach for calibrating LLM confidence with Baysian inference. Our key finding is that using the LLM next token log probability, as proposed by KnowNo [3], is far from enough to adequately calibrate confidence in a real-world setting; both grounding potential actions and applying world knowledge are necessary. Bayesian inference allows us to do just this by conditioning the probability of each possible action on both the grounding in the scene and the available world knowledge. Intuitively, we can think of our Bayesian inference step as integrating the potential for an action to be both possible and safe in the given scene into the confidence measure for each potential action. These steps minimize hallucinations, improve safety, and more accurately calibrate uncertainty. By handling these issues, LBAP can be used to increase the task success rate and minimize the frequency of human intervention. An additional advantage of LBAP is that it does not require extensive training (as suggested in SayCan [1]), which is difficult to do without overfitting for extremely diverse and data-limited tasks like those found in home robotics.
@@
    We introduce LBAP, a novel approach for uncertainty alignment that uses Large Language Models and Bayesian inference to better Align model confidence with task success for high-level robotic planners.
@@
    We specifically propose using Bayesian inference to combine scene grounding and world knowledge with a prior coming from the LLM next token log probabilities.
@@
    We evaluate our approach in both simulation and on hardware in the real world on a ClearPath TurtleBot using a suite of language-instructed manipulation tasks from the KnowNo Simulation and Mobile Manipulator Datasets [3]. We show that LBAP significantly increases the success rate and reduces the amount of help needed as compared to our baselines across different environments and LLMs. Specifically in our real-world testing, we see a decrease in the human help rate of previous methods by over 33% at a success rate of 70%, with a similar decrease at most success rates.
@@
    We also ablate against various terms in the LBAP Bayesian inference, equations 6-8, showing their benefits individually and superior performance when combined.
  5. 5.
    We show that using LBAP on the newest, most powerful language models out-of-the-box (GPT-4) outperforms fine-tuning LBAP and prior art [3] on the best models available for fine-tuning on this task (GPT-3.5).
Figure 2: We present LBAP, a novel approach for uncertainty alignment that utilizes Bayesian inference. After generating a set of multiple-choice candidate actions, each prepended with ‘A’, ‘B’, ‘C’, etc., we find the prior using the LLM log probabilities of each option. Using Bayesian inference we then refine these probabilities by determining the grounding of each option in the scene using the scene context and using world knowledge from LLMs to determine the safety and environmental feasibility of each option. After refining the probabilities of the options and renormalizing, we construct a prediction set of options with a probability above a threshold, $t$. Note that we eliminate option A, ”Put metal bowl in the microwave,” as a valid option as it is dangerous. If the size of the prediction set were greater than one, the agent would ask for help.
@@ II-A Hallucinations and Uncertainty in LLMs
As LLMs have risen to the fore, an increasing amount of work has been aimed at addressing their hallucinations, times when the LLM generates content not grounded in reality. Some works aim to quantify the frequency or severity of hallucinations [4] in a given LLM, while many aim to decrease their frequency [5], or better calibrate uncertainty [9, 10, 11]. Some recent works attempt to handle hallucinations by using conformal prediction-based methods to provide coverage guarantees[12]. Ren et al. [3] introduce KnowNo, which uses conformal prediction to create coverage guarantees for robotics tasks, asking users for help when the agent is unsure. In contrast to these works, we improve the confidence measure and mitigate hallucinations through Bayesian inference, minimizing the amount of human help needed.
II-B Action Feasibility and Robotics
@@
Different from SayCan, we compute all values without the need for training or LLM fine-tuning. We choose to do this as training capable value functions for every possible action in the real world, as SayCan does, is extremely difficult due to a lack of domain data and vast task diversity. We also believe that SayCan’s approach reduces explainability and will hinder the ability to both debug robots in the real-world and explain to their users why they failed. Additionally, we expand upon SayCan in that we incorporate information beyond simply the ability to “successfully execute each skill,” including the safety of executing the action and its conformation to manually set restrictions.
@@ II-C LLMs for Robot Planning and Human-Robot Interaction
A key issue in LLM-based robotics is coming up with a set of candidate actions that respond to the user’s instruction. Prompting LLMs directly for each valid step in a robot plan has proven to perform poorly on uncertain tasks [33]. This becomes a greater issue when we would like to characterize or measure the model’s uncertainty. Furthermore, prior art has shown that the LLM itself is a poor scoring function [33], especially when the plan is expressed in natural language. As such, we choose to build off of prior work, specifically KnowNo, and formulate LLM planning as multiple-choice Q&A (MCQA), a domain LLMs are more familiar with.
III Using Bayesian Inference for Uncertainty Alignment
In this work, we consider the following problem statement: Given a robotic agent and a user-provided instruction, execute the instruction completely and accurately while resolving any ambiguities as needed through dialogue with the user. Minimize the amount of dialogue. Specifically, our robot will begin in a set environment, awaiting a user’s instruction. The user will provide an instruction, which may be under-specified, at which point our method will need to evaluate whether or not it requires additional information from the user. This problem space was initially explored by [3]. For example, if the user asks the robot to put a bowl in the microwave, and there is a plastic and metal bowl on the counter, prior art [3], would need to ask the user for clarification as to which bowl they were referring to. However, our goal for this work is to ensure that the robot knows to use the plastic bowl as putting the metal bowl in the microwave would be dangerous.
@@ III Using Bayesian Inference for Uncertainty Alignment
Following [3], we choose to transform the problem into an multiple choice question answer (MCQA) task, which allows us to define the problem as next-token prediction – aligning well with LLM loss functions and training data. To do this, an LLM is provided a few-shot prompt (provided in supplementary materials) that includes possible next steps in a few example scenarios. The LLM in turn generates a set, $\{y_{i}\}$, of candidate actions that are semantically different for the current scenario. Then, the task of choosing among the candidate actions is formatted as a multiple choice question to the LLM, with each option preceded with ‘A’, ‘B’, ‘C’, ‘D’, or ‘E’ as shown in Figure 2. The probabilities of each possible token (i.e. the multiple choice options ‘A’, ‘B’, ‘C’, …) then serve as normalized scores, or a prior, that can be utilized by uncertainty quantification methods, like LBAP and KnowNo [3]. We utilize these normalized scores as the prior in our Bayesian inference, after which we generate a prediction set of plans from $\{y_{i}\}$. Please see [3] for further rationale on why MCQA should be used for uncertain robot planning.
Bayesian Inference. We choose to use Bayesian inference for uncertainty alignment as it provides a principled framework for updating the confidence estimates from prior knowledge by incorporating multiple sources of additional new information or observed evidence. By applying Bayes’ theorem, $P(A|B)\propto P(B|A)P(A)$, the probability of a hypothesis is adjusted based on both initial expectations and new observations, enabling more accurate confidence calibration for a given task. This approach is particularly beneficial for robotic planning tasks where ambiguity arises from environmental perception, natural language user instructions, or LLM hallucination. In this work specifically, Bayesian inference allows us to condition the likelihood of an action on both scene grounding and world knowledge, better aligning the robot’s confidence measure with the probability of success and thus mitigating hallucinations and reducing the amount of human intervention needed. Bayesian inference provides a lightweight and explainable mechanism for uncertainty alignment unlike methods requiring extensive training.
Method Overview. We present an overview of our method, LBAP, in Figure 2. Given a set of candidate actions ${y_{i}}$ generated by the LLM in the multiple-choice question-answering (MCQA) framework, LBAP aim to estimate the posterior probability
$P(y_{i}|S,W)$ (1)
where $P(y_{i})$ is the probability that an action is the one the user commanded. $S$ is the scene and $W$ is what we are referring to as ‘world knowledge,’ such as knowing that you cannot put a metal bowl in the microwave.
Using Bayes’ theorem (described above), we can express the posterior probability as:
$P(y_{i}|S,W)\propto P(S,W|A)P(y_{i})$ (2)
where $P(y_{i})$ is the prior probability of an action being optimal and $P(S,W|y_{i})$ is the likelihood of observing the given scene and world knowledge if the action is optimal.
Assuming independence between the different sources of evidence given $y_{i}$, we can approximate the likelihood as:
$P(S,W|y_{i})\approx P(S|y_{i})P(W|y_{i})$ (3)
which makes our final equation:
$P(y_{i}|S,W)\propto P(y_{i})P(S|y_{i})P(W|y_{i}).$ (4)
To solve this equation, we break it into its different component parts. As mentioned previously LBAP begins with the MCQA paradigm described above, outputting the model’s uncalibrated confidence score for each generated action option which we take as the prior $P(y_{i})$. We then must compute the other two Bayesian terms, $P(S|y_{i})$ and $P(W|y_{i})$ for each option to calculate the final probability.
@@
LBAP is certain if this set is a singleton, and queries the user for help if not. Calculating an optimal $t$ is the main subject of KnowNo [3], which uses conformal prediction to provide statistical guarantees of a certain success rate. Our work differs significantly from KnowNo as we improve $p(y_{i})$ with Bayesian inference.Our experiments are agnostic to the threshold value used, primarily focusing on improving the relationship between success rate and help rate across all thresholds and subsequent success rates. We ablate against the use of $P(S|y_{i})$ and $P(W|y_{i})$ in our experimentation section.
Grounding the Action in the Scene. We hypothesize that we can estimate $P(S|y_{i})$ by grounding any objects described in $y_{i}$ in the scene. As in some experiments, and in many of KnowNo’s evaluations, we do not have the ability to access the scene directly through perception, we use the textual description of the scene as a fiat. For these scenarios, we use the context of the scene that in theory the robot has already found and provided to the LLM from its perception algorithms. Specifically, to create the prompts for both MCQA and the scoring of the MCQA options, context about the scene was found, we reuse this context. For example, in the simulation environment, there are three blocks and three bowls, one of each is blue, green, and yellow, respectively. The robot’s perception algorithms found these objects and conveyed their existence to the agent as part of its initial prompts. While perception information is provided, this does not always prevent the LLM from hallucinating new objects into existence, like a ‘gold bowl.’ These types of hallucinations become even more prominent when the hallucinated object is in the user command. An example of this can be seen in Figure 1 with the ‘coffee mug’ object.
@@
To calculate $P(S|y_{i})$, the scene context is first parsed into a set of found objects, $\mathcal{O}$. Each proposed action, $y_{i}$, is then itself parsed for any objects, $o^{i}$, present. If all of the objects in $y_{i}$ are in the set of found objects, $a^{i}_{c}=1$. Otherwise, $a^{i}_{c}=\epsilon$, a small probability for rare cases. To continue our example, if $y_{i}$ included a ‘blue bowl’ and a ‘yellow block,’ then $a^{i}_{c}=1$. If $y_{i}$ included a ‘gold bowl,’ then $a^{i}_{c}=\epsilon$. To formalize this:
$P(S|y_{i})=\begin{cases}1&\text{if }\forall o^{i}\in y_{i},\,o^{i}\in\mathcal{O}\\ \epsilon&\text{otherwise}.\end{cases}$ (6)
@@
Grounding in the Scene with Perception. If perception is directly available, and time does not prohibit its use, taking into account any uncertainty in the perception algorithms themselves is valuable. To do this in our experiments, we run an open-vocabulary vision-language model, like CLIP [34] or ViLD [35], on the scene observations with each object, $o^{i}$ (from the option generated by the LLM, $y_{i}$) as input.
@@
Specifically, we take the camera images of the scene, $S$, and feed them through ViLD with the list of objects $\{o_{i}\}$ as input. We then take the probabilities of each object $p(o_{i}|s)$ as outputted by the model and find the average for a given option, $y_{i}$. Formally, we define $P(S|y_{i})$ as
$P(S|y_{i})=\prod_{o_{i}\in y_{i}}p(o_{i}|s),$ (7)
@@
for $n$ objects in $y_{i}$. If the object is localized to the same location as an object in the scene context using IoU (usually an indicator of a hallucination), the metric is lowered to $\epsilon$ to prevent duplicate options of the same object.
@@
Applying World Knowledge. Recent works like [5] and [22] have shown that additional prompts to the LLM requiring it to reason on its prior outputs can isolate harmful hallucinations. We leverage this idea to estimate $P(W|y_{i})$, using the world knowledge in LLMs to determine if the generated action conforms to it. This approach also allows for flexibility, where our prompt or sequence of prompts can account for factors beyond just the grounding of objects in the scene, like the safety of completing a given action. Our prompt, given our real-world evaluation environment, begins with a series of few-shot samples, before serving the following text:
@@
In the prompt, the items inside the brace notation are replaced with their respective values. The header includes the description of items found in the scene from the MCQA generation and scoring prompts and describes the task for the LLM: determining whether the given action is ‘possible and safe’ in the current scene. The header also tells the LLM to reply in either ‘True’ or ‘False’ and nothing else, as well as provides a few examples. Once we receive the response from the LLM, we extract the normalized probability of the ‘True’ token, $p("True"|pr)$ and utilize this as $P(W|y_{i})$ for the option $y_{i}$. Notable is that we believe this approach can be applied across different prompts eliciting different aspects of ‘world knowledge.’ For example, if we have some use case that has business rules such as ‘Do not leave the living room’ we can create a new prompt $pr_{i}$ following the above structure and evaluate that requirement. Creating a set of prompts $PR$ and combining their results gives us:
$P(W|y_{i})=\prod_{pr_{i}\in PR}p("True"|pr_{i}).$ (8)
In the version of LBAP we evaluate in our paper, we only employ the provided above prompt. By evaluating these conditions now instead of in the initial prompt to the LLM ensures that they can be evaluated independently, allowing for greater explainability in our overall system and significantly improved performance.
Figure 3: Comparison of task success rate vs human help rate (left) and average prediction set size (right) in our Simulation experiments. Note LBAP improving upon KnowNo for almost all success rates. Including a reduction in the human help rate of over 70% at a success rate of 60%. No Help and Binary do not create prediction sets and are not included in the plot.
Fine-Tuning. We also hypothesize that utilizing the newest, most powerful, LLM available off-the-shelf will generally outperform fine-tuning the older models available for fine-tuning. We tested this hypothesis by fine-tuning gpt-3.5-turbo-1106 [36], the most powerful model available from OpenAI for fine-tuning at the time of our experiments. For the simulation environment, we fine-tuned on each of the three LLM phases of LAP, the MCQA generation, MCQA scoring, and A-Feasibility scoring.